OpenStudy (anonymous):

rewrite with only sinx and cosx : sin 3x + cos3x

OpenStudy (anonymous):

sin(2x +x)+cos(2x+x)

OpenStudy (anonymous):

2sinxcosxcosx+2cos^2x-1cosx

OpenStudy (mertsj):

sin(3x)= sin(x+2x)=sinxcos2x+cos2xsinx cos(3x)=cos(x+3x)=cosxcos2x-sinxsin2x

OpenStudy (mertsj):

sin2x=2sinxcosx cos2x=cos^2x-sin^2x

OpenStudy (mertsj):

\[sinx(\sin^2x-\cos^2x)+cosx(2sinxcosx)+cosx(\sin^2x-\cos^2x)-sinx(2sinxcosx)\]

OpenStudy (mertsj):

The missing part is: \[-sinx(2sinxcosx)\]

OpenStudy (mertsj):

Simplifies to -sin^3x+cos^3x+3sinxcos^2-3sin^2xcosx